3. (10%) Calculate the leakage current in the nMOS transistor of an inverter with a V = 0.13 volts. Assume V, =0.58 volts, n= 1, y=0, n=0.12, Vod=4 volts, and left = 3.3 nA. Assume it is at room temperature where vi= 0.026 volts. Note that these conditions imply the output of the inverter will be high or at VDD. Ids =
Answer #1

The leakage current in the MOSFET is given given by,

\(I_{l e a k}=I_{o f f} e^{\left(V_{g s}-V_{t}\right) / n v_{T}}\left[1-e^{V_{d d} / v_{T}}\right]\)

\(I_{\text {leak }}=3.3 \times e^{(0.13-0.58) /(0.12 \times 0.026)}\left[1-e^{4 / 0.026}\right] n A\)

\(I_{\text {leak }}=-3.3 \times 3 \times 10^{-63}\left[-6.524 \times 10^{66}\right] n A=61 \mu A\)

Therefore, the leakage current is \(61 \mu A\).

answered by: ANURANJAN SARSAM